A man earns Rs. 4.40 per hour for a 40 hour working week. For overtime done on mondays to fridays, he receives one and half times the normal hourly rate, and for any saturday work, he is paid two times the normal rate. Calculate his wages for 64 hours week of which 4 hours are for work on a saturday.
Answer by kara!(17)

$4.40*40= $176 ( the man will get paid when he will work 40 hours)
64-40=24 hour (worked owertime hours)
from overtime hours of 4hours he worked on Saturday and got paid 2 tomes the normal rate, so 2*$4.40*4 hours=$35.20
on Friday and monday he worked total 20 hours overtime and got paid 1.5 times the normal rate, so 1.5*4.40*20hours=$132
The man earned total of $176+$132+$35.20=$343.20/ weekly